# Nonlinear Schrödinger-Poisson systems in dimension two: the zero mass case

### v1

@@ 4.1. Mountain pass level estimate in the critical case
Therefore, combining (5.11), (5.12), and (5.13) we infer
(5.14) $\begin{split}\text{o}_{n}(1)&=\|\widetilde{u}_{n}-\widetilde{u}\|^{2}_{D^{1,2}}+A_{1}(F(\widetilde{u}_{n}),f(\widetilde{u}_{n})(\widetilde{u}_{n}-\widetilde{u}))\\ &=\|\widetilde{u}_{n}-\widetilde{u}\|^{2}_{D^{1,2}}+A_{1}(F(\widetilde{u}_{n}),\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})(\widetilde{u}_{n}-\widetilde{u}))\\ &\quad+A_{1}\left(F(\widetilde{u}_{n}),\left(\widetilde{f}(\widetilde{u}_{n})-\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})\right)(\widetilde{u}_{n}-\widetilde{u})\right)\end{split}$
@@
one could apply Lemma 3.1(b) and obtain $\|\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})(\widetilde{u}_{n}-\widetilde{u})\|_{\ast,1}\to 0$, which in turn implies that $\|\widetilde{u}_{n}-\widetilde{u}\|_{\ast,p}\to 0$ as $n\to+\infty$ by ($f_{sc}$) or ($f_{c}$). Together with $\widetilde{u}_{n}\to\widetilde{u}$ in $D^{1,2}(\mathbb{R}^{2})$, this would conclude the proof. Since the first two terms in the right-hand side of (5.14) are positive, we are then lead to show that
@@
(5.18) $\begin{split}\bigg{|}\int_{B_{R}(0)}&\!\big{(}\ln(b+|\cdot|)\ast F(\widetilde{u}_{n})\big{)}\left[\left(\widetilde{f}(\widetilde{u}_{n})-\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})\right)(\widetilde{u}_{n}-\widetilde{u})\right]\,{\rm d}x\bigg{|}\\ &\leq\|F(\widetilde{u}_{n})\|_{\ast,1}\int_{\mathbb{R}^{2}}\left|\widetilde{f}(\widetilde{u}_{n})-\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})\right||\widetilde{u}_{n}-\widetilde{u}|\\ &\quad+\|F(\widetilde{u}_{n})\|_{1}\int_{\mathbb{R}^{2}}\ln(b+|x|)\left|\widetilde{f}(\widetilde{u}_{n})-\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})\right||\widetilde{u}_{n}-\widetilde{u}|\,{\rm d}x\,.\end{split}$
@@
(5.19) $\begin{split}\int_{\mathbb{R}^{2}}f(\widetilde{u}_{n})|\widetilde{u}_{n}-\widetilde{u}|&\lesssim\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}|^{p-1}|\widetilde{u}_{n}-\widetilde{u}|+\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}|^{p-1}|\widetilde{u}_{n}-\widetilde{u}|\,{\rm e}^{\alpha\tilde{u_{n}}^{2}}\\ &\leq\|\widetilde{u}_{n}\|_{p}^{p-1}\|\widetilde{u}_{n}-\widetilde{u}\|_{p}+\left(\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}|^{p}\,{\rm e}^{\alpha\frac{p}{p-1}\tilde{u_{n}}^{2}}\right)^{\frac{p-1}{p}}\|\widetilde{u}_{n}-\widetilde{u}\|_{p}\\ &\lesssim\|\widetilde{u}_{n}\|_{p}^{p-1}\|\widetilde{u}_{n}-\widetilde{u}\|_{p}=\text{o}_{n}(1)\,,\end{split}$
by Theorem 2.2 with $\alpha<4\pi\frac{p-1}{p}\left(\sup_{n}\|\nabla\widetilde{u}_{n}\|_{2}\right)^{-2}$. By similar estimate, one may also show that $\int_{\mathbb{R}^{2}}f(\widetilde{u}_{n})|\widetilde{u}_{n}-\widetilde{u}|=\text{o}_{n}(1)$. On the other hand, if $f$ is critical, namely ($f_{c}$) holds, defining now $\widetilde{v}_{n}=G(\widetilde{u}_{n})$ with $G$ as in (4.35) and arguing as in Lemma 4.6, one finds
(5.20) $\begin{split}\int_{\mathbb{R}^{2}}f(\widetilde{u}_{n})|\widetilde{u}_{n}-\widetilde{u}|&\lesssim\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}|^{p-1}|\widetilde{u}_{n}-\widetilde{u}|+\int_{\mathbb{R}^{2}}\left(t_{\varepsilon}+\frac{\widetilde{v}_{n}}{1-\varepsilon}\right)^{r-1}|\widetilde{u}_{n}-\widetilde{u}|{\rm e}^{\alpha\left(t_{\varepsilon}+\frac{\widetilde{v}_{n}}{1-\varepsilon}\right)^{2}}\\ &\lesssim\|\widetilde{u}_{n}\|_{p}^{p-1}\|\widetilde{u}_{n}-\widetilde{u}\|_{p}+\!\left(\!\int_{\mathbb{R}^{2}}\!\left(t_{\varepsilon}+\frac{\widetilde{v}_{n}}{1-\varepsilon}\right)^{(r-1)\sigma^{\prime}}\!\!{\rm e}^{\alpha\sigma^{\prime}\left(t_{\varepsilon}+\frac{\widetilde{v}_{n}}{1-\varepsilon}\right)^{2}}\right)^{\frac{1}{\sigma^{\prime}}}\!\|\widetilde{u}_{n}-\widetilde{u}\|_{\sigma}\\ &=\text{o}_{n}(1)\,,\end{split}$
@@
(5.21) $\begin{split}\int_{\mathbb{R}^{2}}\big{|}\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})\big{|}&|\widetilde{u}_{n}-\widetilde{u}|\lesssim\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}-\widetilde{u}|^{p}+\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}-\widetilde{u}|^{r}{\rm e}^{\alpha|\widetilde{u}_{n}-\widetilde{u}|^{2}}\\ &\leq\text{o}_{n}(1)+\left(\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}-\widetilde{u}|^{r}{\rm e}^{\varepsilon\alpha s\,\tilde{u_{n}}^{2}}\right)^{\frac{1}{s}}\left(\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}-\widetilde{u}|^{r}{\rm e}^{C_{\varepsilon}\alpha s^{\prime}\,\tilde{u}^{2}}\right)^{\frac{1}{s^{\prime}}}\\ &\leq\text{o}_{n}(1)+\left(C\int_{\{\widetilde{u}_{n}\leq 1\}}|\widetilde{u}_{n}-\widetilde{u}|^{r}+\int_{\{\widetilde{u}_{n}>1\}}|\widetilde{u}_{n}-\widetilde{u}|^{r}|\widetilde{u}_{n}|^{\kappa}{\rm e}^{\varepsilon\alpha s\,\tilde{u_{n}}^{2}}\right)^{\frac{1}{s}}\\ &\quad\times\left(C\int_{\{\widetilde{u}\leq 1\}}|\widetilde{u}_{n}-\widetilde{u}|^{r}+\int_{\{\widetilde{u}>1\}}|\widetilde{u}_{n}-\widetilde{u}|^{r}|\widetilde{u}|^{\kappa}{\rm e}^{C_{\varepsilon}\alpha s^{\prime}\,\tilde{u}^{2}}\right)^{\frac{1}{s^{\prime}}}\\ &\leq\text{o}_{n}(1)+\left(\text{o}_{n}(1)+\|\widetilde{u}_{n}-\widetilde{u}\|_{rq^{\prime}}^{r/s}\left(\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}|^{\kappa q}{\rm e}^{\varepsilon\alpha sq\,\tilde{u_{n}}^{2}}\right)^{\frac{1}{sq}}\right)\\ &\quad\times\left(\text{o}_{n}(1)+\|\widetilde{u}_{n}-\widetilde{u}\|_{rq^{\prime}}^{r/s^{\prime}}\left(\int_{\mathbb{R}^{2}}|\widetilde{u}|^{\kappa q}{\rm e}^{C_{\varepsilon}\alpha s^{\prime}\,{\tilde{u}^{2}}}\right)^{\frac{1}{s^{\prime}q}}\right).\end{split}$
Choosing now the parameters $\alpha>\alpha_{0}$ close to $\alpha_{0}$ and $\kappa,r,q,s>1$ so that $rq^{\prime}>p$, $\kappa q>p$ and $\varepsilon>0$ such that $\varepsilon\alpha_{0}qs\|\nabla\widetilde{u}_{n}^{2}\|<4\pi$, which is possible since $\{\widetilde{u}_{n}\}_{n}$ is bounded in $D^{1,2}(\mathbb{R}^{2})$, then one may apply Theorem 2.2 and, combining (5.21) with (5.19) or (5.20), get
(5.22) $\int_{\mathbb{R}^{2}}\big{|}\widetilde{f}(\widetilde{u}_{n})-\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})\big{|}|\widetilde{u}_{n}-\widetilde{u}|\lesssim\text{o}_{n}(1)+\left(\text{o}_{n}(1)+\text{o}_{n}(1)C\|\widetilde{u}_{n}\|_{p}^{p/sq}\right)\left(\text{o}_{n}(1)+\text{o}_{n}(1)C\|\widetilde{u}\|_{p}^{p/sq}\right)=\text{o}_{n}(1)\,,$
since $\{\widetilde{u}_{n}\}_{n}$ is also bounded in $L^{p}(\mathbb{R}^{2})$. It remains to prove that also the last term in (5.18) vanishes. To this aim, since $f\in C^{1}(\mathbb{R})$ and so is $\widetilde{f}$ (cf. (5.15)), by Lagrange’s theorem,
(5.23) $\widetilde{f}(\widetilde{u}_{n})-\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})=\widetilde{f}^{\prime}(w_{n})\widetilde{u}=f(|w_{n}|)\widetilde{u},\quad\ \mbox{where}\quad w_{n}:=\theta_{n}\widetilde{u}_{n}+(1-\theta_{n})\widetilde{u}$
with $\theta_{n}:\mathbb{R}^{2}\to[0,1]$. Therefore, splitting $\mathbb{R}^{2}=B_{R}(0)\cup B_{R}(0)^{c}$ for a fixed $R>0$, we obtain
(5.24) $\begin{split}&\int_{\mathbb{R}^{2}}\ln(b+|x|)\left|\widetilde{f}(\widetilde{u}_{n})-\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})\right||\widetilde{u}_{n}-\widetilde{u}|\,{\rm d}x\\ &\ \ \leq\ln(b+R)\!\int_{B_{R}(0)}\!\big{|}\widetilde{f}(\widetilde{u}_{n})-\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})\big{|}|\widetilde{u}_{n}-\widetilde{u}|+\!\int_{B_{R}(0)^{c}}\!\!\ln(b+|x|)|f^{\prime}(|w_{n}|)|\,\widetilde{u}\,|\widetilde{u}_{n}-\widetilde{u}|\,{\rm d}x\\ &\ \ \leq\text{o}_{n}(1)+\left(\int_{\mathbb{R}^{2}}\ln(b+|x|)|f^{\prime}(|w_{n}|)|^{\frac{p}{p-2}}\,{\rm d}x\right)^{\frac{p-2}{p}}\!\left(\int_{B_{R}(0)^{c}}\ln(b+|x|)|\widetilde{u}|^{p}\,{\rm d}x\right)^{\frac{1}{p}}\!\left(\|\widetilde{u}_{n}\|_{\ast,p}+\|\widetilde{u}\|_{\ast,p}\right)\end{split}$
by (5.22). Since $\|\widetilde{u}\|_{*,p}<+\infty$, for any chosen $\delta>0$ one may find $R>0$ large enough such that (by Lebesgue’s dominated convergence)
(5.25) $\int_{B_{R}(0)^{c}}\ln(b+|x|)|\widetilde{u}|^{p}\,{\rm d}x<\delta\,.$
Recalling (5.7), we are then left to show that
(5.26) $\int_{\mathbb{R}^{2}}\ln(b+|x|)|f^{\prime}(|w_{n}|)|^{\frac{p}{p-2}}\,{\rm d}x\leq C$
for all $n$. By exploiting (5.23), ($f_{5}$), again the $\varepsilon$-Young inequality, (5.7), and $u\in X$, one may estimate it as
(5.27) $\begin{split}\int_{\mathbb{R}^{2}}\ln&(b+|x|)|f^{\prime}(|w_{n}|)|^{\frac{p}{p-2}}\,{\rm d}x\\ &\lesssim\int_{\mathbb{R}^{2}}\ln(b+|x|)(|\widetilde{u}_{n}|^{p}+|\widetilde{u}|^{p})\,{\rm d}x+\int_{\mathbb{R}^{2}}\ln(b+|x|)(\widetilde{u}_{n}+\widetilde{u})^{\frac{pr}{p-2}}\,{\rm e}^{\alpha^{\prime}\frac{p}{p-2}(\tilde{u_{n}}+\tilde{u})^{2}}\,{\rm d}x\\ &\lesssim C+\left(\int_{\mathbb{R}^{2}}\ln(b+|x|)\widetilde{u}_{n}^{p}\,{\rm e}^{\varepsilon\alpha^{\prime}\frac{pq}{p-2}\tilde{u_{n}}^{2}}\,{\rm d}x\right)^{\frac{1}{q}}\left(\int_{\mathbb{R}^{2}}\ln(b+|x|)\widetilde{u}_{n}^{p}\,{\rm e}^{C_{\varepsilon}\alpha^{\prime}\frac{pq^{\prime}}{p-2}\tilde{u}^{2}}\,{\rm d}x\right)^{\frac{1}{q^{\prime}}}\\ &\quad+\left(\int_{\mathbb{R}^{2}}\ln(b+|x|)\widetilde{u}^{p}\,{\rm e}^{\varepsilon\alpha^{\prime}\frac{pq}{p-2}\tilde{u_{n}}^{2}}\,{\rm d}x\right)^{\frac{1}{q}}\left(\int_{\mathbb{R}^{2}}\ln(b+|x|)\widetilde{u}^{p}\,{\rm e}^{C_{\varepsilon}\alpha^{\prime}\frac{pq^{\prime}}{p-2}\tilde{u}^{2}}\,{\rm d}x\right)^{\frac{1}{q^{\prime}}}\end{split}$
by choosing $r=p-2>0$, and we estimate the four terms separately. The first term is easily bounded independently of $n$ if one chooses $\varepsilon$ small enough: in fact, since $\sup_{n\in\mathbb{N}}\|\widetilde{u}_{n}\|\leq\overline{C}$, one has
@@
(5.28) $\int_{\mathbb{R}^{2}}\ln(b+|x|)\widetilde{u}_{n}^{p}\,{\rm e}^{\varepsilon\alpha^{\prime}\frac{pq}{p-2}\tilde{u_{n}}^{2}}\,{\rm d}x=\int_{\mathbb{R}^{2}}\ln(b+|x|)\widetilde{u}_{n}^{p}\,{\rm e}^{\varepsilon\alpha^{\prime}\frac{pq}{p-2}\overline{C}\left(\frac{\tilde{u_{n}}}{\|\tilde{u_{n}}\|}\right)^{2}}\,{\rm d}x\leq C\,,$
by choosing $\varepsilon$ so small that $\varepsilon\alpha^{\prime}\frac{pq}{p-2}\overline{C}\leq(4\pi p)^{-1/2}$ by Theorem 2.3. In a similar way we may prove that the last term in (5.27), which is independent of $n$, is finite. Indeed,
@@
(5.29) $\int_{\mathbb{R}^{2}}\ln(b+|x|)\widetilde{u}^{p}\,{\rm e}^{C_{\varepsilon}\alpha^{\prime}\frac{pq^{\prime}}{p-2}\tilde{u}^{2}}\,{\rm d}x=\int_{\mathbb{R}^{2}}\ln(b+|x|)\widetilde{u}^{p}\,{\rm e}^{C_{\varepsilon}\alpha^{\prime}\frac{pq^{\prime}}{p-2}\|\tilde{u}\|^{2}\left(\frac{\tilde{u}}{\|\tilde{u}\|}\right)^{2}}\,{\rm d}x<+\infty$
for all $\varepsilon>0$, again by Theorem 2.3. Concerning now the second term, which mixes $\widetilde{u}_{n}$ and $\widetilde{u}$, one may estimate as follows:

### v2

@@ 4.1. Mountain pass level estimate in the critical case
Indeed, if ($f_{sc}$) is assumed, it is easy to deal with both terms and prove that the first is uniformly bounded and that the second converges to 00, with analogous estimates as done e.g. in (5.3). On the other hand, if ($f_{c}$) holds, $\|F(\widetilde{u}_{n})\|_{\gamma}\leq C$ by Lemma 4.7, while we can estimate the second term as follows, by defining $v:=G(\widetilde{u}_{n})$ and by means of Lemma 4.6:
$\int_{\mathbb{R}^{2}}\left|f(\widetilde{u}_{n})(\widetilde{u}_{n}-\widetilde{u})\right|^{\gamma}$ $\leq C\|\widetilde{u}_{n}\|_{p\gamma}^{(p-1)\gamma}\|\widetilde{u}_{n}-\widetilde{u}\|_{p\gamma}^{\gamma}+C_{\varepsilon}\|\widetilde{u}_{n}-\widetilde{u}\|_{\sigma^{\prime}\gamma}^{\gamma}$
×$\quad\times\|\nabla v_{n}\|_{2}^{r\gamma}\left(\int_{\mathbb{R}^{2}}\left(\frac{|v_{n}|}{\|\nabla v_{n}\|_{2}}\right)^{r\gamma\sigma}{\rm e}^{\gamma\alpha\sigma\left(\frac{1+\varepsilon}{1-\varepsilon}\right)^{2}\|\nabla v_{n}\|_{2}^{2}\left(\frac{v_{n}}{\|\nabla v_{n}\|_{2}}\right)^{2}}\right)^{\frac{1}{\sigma}}.$
The first term tends to 00 as $n\to+\infty$, since $\{\widetilde{u}_{n}\}_{n}$ is bounded in $X$ and by the continuous and compact embeddings in Lebesgue spaces provided by Proposition 2.1, since $\gamma>1$. By choosing $\gamma>1$ and $\sigma>1$ both close to $1$, $\alpha>\alpha_{0}$ close to $\alpha_{0}$, $\varepsilon>0$ small, and $r=p$, by (4.33) and Theorem 2.2, this yields
$\int_{\mathbb{R}^{2}}\left|f(\widetilde{u}_{n})(\widetilde{u}_{n}-\widetilde{u})\right|^{\gamma}$ $\lesssim\text{o}_{n}(1)+\|\widetilde{u}_{n}-\widetilde{u}\|_{\sigma^{\prime}\gamma}^{\gamma}\|\nabla v_{n}\|_{2}^{p\left(\gamma-\frac{1}{\sigma}\right)}\|v_{n}\|_{p}^{\frac{p}{\sigma}}$
$\lesssim\text{o}_{n}(1)+\left(\frac{4\pi}{\alpha_{0}}\right)^{\frac{p}{2}\left(\gamma-\frac{1}{\sigma}\right)}\|\widetilde{u}_{n}-\widetilde{u}\|_{\sigma^{\prime}\gamma}^{\gamma}\|\widetilde{u}_{n}\|_{p}^{\frac{p}{\sigma}}=\text{o}_{n}(1)\,,$
since $\gamma-\frac{1}{\sigma}>0$ and again by Proposition 2.1. Therefore, combining (5.11), (5.12), and (5.13) we infer
(5.14) $\text{o}_{n}(1)$ $=\|\widetilde{u}_{n}-\widetilde{u}\|^{2}_{D^{1,2}}+A_{1}(F(\widetilde{u}_{n}),f(\widetilde{u}_{n})(\widetilde{u}_{n}-\widetilde{u}))$
$=\|\widetilde{u}_{n}-\widetilde{u}\|^{2}_{D^{1,2}}+A_{1}(F(\widetilde{u}_{n}),\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})(\widetilde{u}_{n}-\widetilde{u}))$
$\quad+A_{1}\left(F(\widetilde{u}_{n}),\left(\widetilde{f}(\widetilde{u}_{n})-\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})\right)(\widetilde{u}_{n}-\widetilde{u})\right)$
@@
one could apply Lemma 3.1(b) and obtain $\|\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})(\widetilde{u}_{n}-\widetilde{u})\|_{\ast,1}\to 0$, which in turn implies that $\|\widetilde{u}_{n}-\widetilde{u}\|_{\ast,p}\to 0$ as $n\to+\infty$ by ($f_{sc}$) or ($f_{c}$). Together with $\widetilde{u}_{n}\to\widetilde{u}$ in $D^{1,2}(\mathbb{R}^{2})$, this would imply that $\widetilde{u}_{n}\to\widetilde{u}$ in $X$, which concludes the proof since $\{\widetilde{u}_{n}\}_{n}$ is a Cerami sequence and $\mathcal{I}$ is of class $C^{1}$. Since the first two terms in the right-hand side of (5.14) are positive, we are then lead to show that
@@
(5.18) $\bigg{|}\int_{B_{R}}$ $\!\big{(}\ln(b+|\cdot|)\ast F(\widetilde{u}_{n})\big{)}\left[\left(\widetilde{f}(\widetilde{u}_{n})-\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})\right)(\widetilde{u}_{n}-\widetilde{u})\right]\bigg{|}$
$\leq\|F(\widetilde{u}_{n})\|_{\ast,1}\int_{\mathbb{R}^{2}}\left|\widetilde{f}(\widetilde{u}_{n})-\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})\right||\widetilde{u}_{n}-\widetilde{u}|$
$\quad+\|F(\widetilde{u}_{n})\|_{1}\int_{\mathbb{R}^{2}}\ln(b+|x|)\left|\widetilde{f}(\widetilde{u}_{n})-\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})\right||\widetilde{u}_{n}-\widetilde{u}|\,.$
@@
$\int_{\mathbb{R}^{2}}f(\widetilde{u}_{n})|\widetilde{u}_{n}-\widetilde{u}|$ $\lesssim\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}|^{p-1}|\widetilde{u}_{n}-\widetilde{u}|+\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}|^{p-1}{\rm e}^{\alpha\widetilde{u}_{n}^{2}}\,|\widetilde{u}_{n}-\widetilde{u}|$
(5.19) $\leq\|\widetilde{u}_{n}\|_{p}^{p-1}\|\widetilde{u}_{n}-\widetilde{u}\|_{p}+\left(\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}|^{p}\,{\rm e}^{\alpha\frac{p}{p-1}\widetilde{u}_{n}^{2}}\right)^{\frac{p-1}{p}}\!\|\widetilde{u}_{n}-\widetilde{u}\|_{p}$
$\lesssim\|\widetilde{u}_{n}\|_{p}^{p-1}\|\widetilde{u}_{n}-\widetilde{u}\|_{p}=\text{o}_{n}(1)\,,$
by Theorem 2.2 taking $\alpha<4\pi\frac{p-1}{p}\left(\sup_{n}\|\nabla\widetilde{u}_{n}\|_{2}\right)^{-2}$. By a similar estimate, one may also show that $\int_{\mathbb{R}^{2}}\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})|\widetilde{u}_{n}-\widetilde{u}|=\text{o}_{n}(1)$.
On the other hand, if $f$ is of critical growth, namely ($f_{c}$) holds, defining $\widetilde{v}_{n}=G(\widetilde{u}_{n})$, by Lemma 4.6 one finds
(5.20) $\int_{\mathbb{R}^{2}}f(\widetilde{u}_{n})|\widetilde{u}_{n}-\widetilde{u}|$ $\lesssim\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}|^{p-1}|\widetilde{u}_{n}-\widetilde{u}|+\int_{\mathbb{R}^{2}}\left(t_{\varepsilon}+\tfrac{\widetilde{v}_{n}}{1-\varepsilon}\right)^{r-1}{\rm e}^{\alpha\left(t_{\varepsilon}+\frac{\widetilde{v}_{n}}{1-\varepsilon}\right)^{2}}|\widetilde{u}_{n}-\widetilde{u}|$
$\lesssim\|\widetilde{u}_{n}\|_{p}^{p-1}\|\widetilde{u}_{n}-\widetilde{u}\|_{p}+\!\left(\!\int_{\mathbb{R}^{2}}\!\left(t_{\varepsilon}+\tfrac{\widetilde{v}_{n}}{1-\varepsilon}\right)^{(r-1)\sigma^{\prime}}\!\!{\rm e}^{\alpha\sigma^{\prime}\left(t_{\varepsilon}+\frac{\widetilde{v}_{n}}{1-\varepsilon}\right)^{2}}\right)^{\frac{1}{\sigma^{\prime}}}\!\|\widetilde{u}_{n}-\widetilde{u}\|_{\sigma}$
$=\text{o}_{n}(1)\,,$
@@
$\int_{\mathbb{R}^{2}}\big{|}\widetilde{f}(\widetilde{u}_{n}-\widetilde{u})\big{|}|\widetilde{u}_{n}-\widetilde{u}|$ $\lesssim\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}-\widetilde{u}|^{p}+\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}-\widetilde{u}|^{r}{\rm e}^{\alpha|\widetilde{u}_{n}-\widetilde{u}|^{2}}$
@@
$\leq\text{o}_{n}(1)+\left(\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}-\widetilde{u}|^{r}{\rm e}^{(1+\varepsilon)\alpha s\,\widetilde{u}_{n}^{2}}\right)^{\frac{1}{s}}\left(\int_{\mathbb{R}^{2}}|\widetilde{u}_{n}-\widetilde{u}|^{r}{\rm e}^{C_{\varepsilon}\alpha s^{\prime}\,\widetilde{u}^{2}}\right)^{\frac{1}{s^{\prime}}}$
@@
(5.21) $\leq\text{o}_{n}(1)+\left(C\!\int_{\{\widetilde{u}_{n}\leq 1\}}\!|\widetilde{u}_{n}-\widetilde{u}|^{r}+\int_{\{\widetilde{u}_{n}>1\}}\!|\widetilde{u}_{n}-\widetilde{u}|^{r}|\widetilde{u}_{n}|^{\kappa}{\rm e}^{(1+\varepsilon)\alpha s\,\widetilde{u}_{n}^{2}}\right)^{\frac{1}{s}}$